A pool measuring 60ft by 26ft needs 1ft squares tiles around the deck. How many tiles are needed?
Answer by Alan3354(69443)

A pool measuring 60ft by 26ft needs 1ft squares tiles around the deck. How many tiles are needed?
-------------
The perimeter is 2W + 2L = 2*60 + 2*26 = 172 feet
---> 172 tiles + 1 for each corner
= 176 tiles
========================
Is there some reason you could not figure that out?